Home | Contact Us | Submit Troubleshooting Info | Search | Help


    Output | Software Solutions | Proofing | Color, Screening and Calibration | Scanning | Digital Photography | Consumables | POS


Using Head Firmware 1.01F with Version 3.50.05/3.50.07

The following problem can occur with certain TH2 heads that are currently in the field or in the integration line:
The data line that carries packets that include the Write Laser temperature set point from VER D T2HC to Ver D WLPB are susceptible to noise. This can cause the laser temperature to be set to an erroneous value.
The problem can occur when using firmware version 1.01e or earlier. The MCE firmware does not notify the head that the shutter will open in order for the head to stop monitoring laser power. Therefore, the head powermonitor mode must be turned off. This mode did not exist in firmware versions prior to, and including, the 1.01 series. In upgrades to series 1.01, Off is the default.

Head firmware 1.01F solves this problem. Since the firmware was checked only as part of the 3.51.02 Quantum package, in order to use these (known) defective heads with Quantum versions prior to 3.51.02 (3.50.05 and 3.50.07), load only the new head firmware, change a certain head parameter (only once) and continue to use the regular version safely.

Loading Head Firmware 1.01F (attached below in zip format):

  1. Copy the file th2h101f.bin to the machine's C:\th2\tools directory.
  2. Close the Lotem application.
  3. Open SPlot.
  4. Type: QUANTUM>> set head type 3
  5. Type: QUANTUM>> head program
  6. Type: QUANTUM>> head boot
    and wait for percentage completion.
  7. Close Splot.
  8. Open a DOS prompt.
  9. Type: C:\th2\tools\ tehload p=8 th2h101f.bin
    and wait for percentage completion.
  10. Open SPlot.
  11. Type: QUANTUM>> reset

Turning Head Power Monitor Mode Off:

  1. Open SPlot .
  2. Type: QUANTUM>> th2 "eng"
  3. Type: QUANTUM>> th2 "set sys powermonitor false"
  4. Type: QUANTUM>> th2 "set sys" and verify that the powermonitor line appears as follows:
    [03510] SYS powermonitor False True

    Head S/N DU Customer's name Quantum S/N Comments
    T00375-1       No records
    T00384-1 CAM Spares    
    T00388-1 CEU KRAFT DRUCK - GER 0210-103446 Prinergy
    T00389-1 CEU Spares    
    T00390-1 CEU FRANZ SPIEGEL BUCH 0210-102535 Prinergy
    T00391-1 CEU Unwin Brothers - UK 0210-102534  
    T00402-1 CEU MCSI 0211-104006 Brisque
    T00403-1 CAM Spares    
    T00404-2 CIL Production line    
    T00405-1 CEU St. Ives 0211-103447 Brisque
    T00406-1 CIL On the way from CCA    
    T00412-1 CAP Printing Express Ltd. HK 0211-104005 Brisque
    T00416-1 CIL Production line    
    T00419-1 CAP Spares    
 

    Copyright © 2005 Creo Inc. All rights reserved. Confidential. Back to top.